OpenStudy (anonymous):

solve each system using the elimination method. x+3y=1 -x+5y=-9

OpenStudy (anonymous):

x+3y=1 -x+5y=-9 ---------- 8y = -8 y = -1 Then plug -1 in the first equation. x + 3(-1) = 1 x = 4 So the answer is y = -1, x = 4.

OpenStudy (anonymous):

I can confirm, that ashrafh is in fact right.

OpenStudy (anonymous):

@ashrafh 5x-2y=-6 4x+5y=-18